# Formalizing the Safety, Security, and Functional Properties of Agentic AI Systems

### v1

@@ Abstract
Agentic AI systems, which leverage multiple autonomous agents and Large Language Models (LLMs), are increasingly used to address complex, multi-step tasks. The safety, security, and functionality of these systems are critical, especially in high-stakes applications. However, the current ecosystem of inter-agent communication is fragmented, with protocols such as the Model Context Protocol (MCP) for tool access and the Agent-to-Agent (A2A) protocol for coordination being analyzed in isolation. This fragmentation creates a semantic gap that prevents the rigorous analysis of system properties and introduces risks such as architectural misalignment and exploitable coordination issues. To address these challenges, we introduce a modeling framework for agentic AI systems composed of two foundational models. The first, the host agent model, formalizes the top-level entity that interacts with the user, decomposes tasks, and orchestrates their execution by leveraging external agents and tools. The second, the task lifecycle model, details the states and transitions of individual sub-tasks from creation to completion, providing a fine-grained view of task management and error handling. Together, these models provide a unified semantic framework for reasoning about the behavior of multi-AI agent systems. Grounded in this framework, we define 17 properties for the host agent and 14 for the task lifecycle, categorized into liveness, safety, completeness, and fairness. Expressed in temporal logic, these properties enable formal verification of system behavior, detection of coordination edge cases, and prevention of deadlocks and security vulnerabilities. Through this effort, we introduce the first rigorously grounded, domain-agnostic framework for the systematic analysis, design, and deployment of correct, reliable, and robust agentic AI systems.
I Introduction
@@
Agentic AI systems are becoming central to the creation of advanced AI. These systems address complex, multi-step tasks that surpass the capabilities of any single agent by utilizing multiple autonomous agents. Agents in these systems employ Large Language Models (LLMs) as their primary reasoning engine, which enables them to perform task planning, delegation, and complex decision-making [1, 2, 3]. These heterogeneous agents (e.g., agents based on language, vision, robotics, and symbolic planning) dynamically communicate, delegate, and collaborate to complete predefined tasks.
@@
To illustrate, we consider an automated financial planner. A host agent accepts a natural language request from a user to “create a budget and invest $\mathdollar 1,000$”. It decomposes this into subtasks: a data agent gathers market data, a planning agent calculates risk, and a transaction agent executes the final investment. These specialized agents interact and exchange data to accomplish the multi-step task.
The operation of such systems relies on standardized communication protocols, such as Model Context Protocol (MCP) [4], which addresses vertical tool access for a single agent, and Agent-to-Agent (A2A) [5] protocol, which addresses horizontal agent coordination via inter-agent delegation. The MCP allows an agent to discover, authenticate with, and invoke external tools and data sources in a structured manner. The A2A protocol provides mechanisms for agents to discover peer capabilities, negotiate task assignments, and coordinate execution across distributed environments.
@@
Turning back to the financial planner example, the host agent must leverage both protocols. It uses the A2A protocol to delegate the subtask of gathering market data from the data agent to the planning agent (horizontal agent coordination). Conversely, the transaction agent relies on the MCP protocol to invoke an external banking API tool to execute the final investment (vertical tool access). The ability of the host agent to manage and sequence both A2A and MCP interactions across various agents enables the multi-step tasks to succeed.
@@
The formal or empirical assessment of safety, security, and functionality of multi-AI agent systems is increasingly important as these systems transition to open-ended, high-stakes applications [6, 7]. Here, safety involves preventing the agent’s actions from causing unintended or harmful physical or real-world consequences, such as an incorrect trade that causes a substantial loss in the financial planner example. Functionality ensures that the system accurately and reliably completes the specified task, such as verifying that the financial planner calculates risk correctly and executes the investment as requested. Safety and functionality issues might happen naturally, e.g., due to design flaws or code errors, but security considers adversarial manipulation and unauthorized access that can violate such safety and functionality requirements. For instance, a malicious external agent could use a prompt injection attack [8] to force the transaction agent into transferring funds to an unauthorized account, a clear violation of both security and safety policies.
@@
Recent work has directly or indirectly started studying these issues by examining agents in specific settings, such as adversarial agent behavior [9, 10, 11] and protocol misuse in task delegation [12]. Other lines of research focus on the compromise or malfunction of a single agent that propagates errors, escalates failures, or exposes vulnerabilities across the entire system [13, 14, 15, 16]. Further studies also examine the risk of using combinations of safe models for misuse [17], the security challenges inherent in tool usage [18], and ensuring secure delegation and preventing protocol misuse in adversarial environments [19, 20].
@@
The current ecosystem of inter-agent communication and coordination is fragmented and lacks a cohesive formal foundation. Existing protocols (e.g., MCP and A2A) are analyzed in isolation, which creates a semantic gap that impedes rigorous reasoning about safety, security, and functional properties when both are used simultaneously in complex workflows. This absence of a unified framework, a need that is also recognized by multi-stakeholder initiatives [21], leaves deployments vulnerable to emergent coordination failures, including deadlocks and privilege escalation, that cannot be adequately verified or mitigated. Achieving verifiable correctness in these systems requires a rigorous, unified modeling framework that can capture the execution behavior of diverse agents across heterogeneous protocols.
Our work addresses these challenges by introducing a modeling framework for agentic AI systems that consists of two foundational models. These models represent the conceptual representations of the system’s architecture and behavior:
@@
    Host Agent (HA) Model: This model formalizes the top-level entity that interacts with the user. The HA is responsible for accepting user tasks, decomposing them into structured subtasks, and orchestrating execution by leveraging external entities (agents via A2A and tools via MCP). It acts as both a controller and a monitor, ensuring safe delegation and state consistency throughout the task lifecycle.
@@
    Task Lifecycle Model: This model formalizes the dynamic structure of task management, detailing the states and transitions an individual sub-task undergoes from its origin to its completion or failure. It provides the fine-grained execution logic necessary for dependency management and resilient error handling.
These models are the foundational structure that enables formal verification. Specifically, they capture the discrete states and transitions of the entire coordination process and provide the necessary semantic framework (state space) to express the formal safety, security, and functional requirements. This unified perspective is required to rigorously reason about execution behavior, uncover coordination edge cases, and verify properties critical to real-world deployments.
We categorize the formal properties defined within this framework as liveness, safety, completeness, and fairness, which are essential for verifiable system assurance. For instance, a safety property guarantees that the transaction agent will not execute an investment before the planning agent has successfully calculated the risk, thus it prevents inconsistent state or premature execution. Conversely, a liveness property ensures that once a user submits a request, a final response is eventually returned. This prevents the financial planning task from entering a permanent deadlock or starvation state.
In summary, we make the following contributions:
  • We introduce a host agent model, which provides a rigorous and domain-agnostic foundation for the systematic analysis of multi-AI agent systems. This model formalizes the top-level agent that interacts with the user, decomposes tasks, and orchestrates their execution by leveraging external agents and tools.
@@
II Background
II-A Agentic AI systems
AI agents are autonomous computational entities that perceive their environment, reason about goals, and take actions to achieve desired outcomes [22]. Unlike traditional pre-programmed software, AI agents exhibit adaptive behavior, making decisions based on incomplete or uncertain information. The field progressed from early symbolic approaches [23] to systems that use LLMs as their reasoning engines [24, 25].
LLM-based agents show potential for human-level intelligence through access to vast web knowledge [26]. These systems process natural language, plan multi-step tasks, use tools, generate code, analyze data, and perform complex reasoning [27]. However, single-agent approaches exhibit limitations when tasks become sophisticated and multi-faceted. Complex scenarios require specialized expertise across multiple domains, concurrent execution of interdependent subtasks, and coordination mechanisms that an individual agent cannot manage [28, 29, 1].
Agentic AI systems represent a paradigm shift toward collaborative networks of specialized agents [30, 31]. In these systems, autonomous agents interact, communicate, and coordinate their actions to solve problems beyond individual scope [32]. These systems leverage distributed intelligence to decompose tasks and allocate them to agents based on specialized capabilities and current availability.
II-B Agent Communication and Coordination Frameworks
@@
Agentic AI systems require standardized frameworks for individual agent capabilities and multi-agent coordination. These frameworks comprise two complementary aspects: protocols for agents to interact with external resources and tools, and protocols for inter-agent communication. For individual agents, the challenge is ensuring secure, reliable access to external computational resources and tools. Agents must discover capabilities, authenticate services, validate parameters, and handle errors robustly. For multi-agent coordination, protocols define how agents discover peer capabilities, negotiate task assignments, exchange information, and maintain consistency across distributed execution contexts.
The MCP [4] and A2A [5] protocols illustrate the integration challenges that arise from combining individual agent capabilities with multi-agent coordination.
Model Context Protocol (MCP): Agent-Tool Integration. MCP addresses structured tool access for individual Agentic AI [4]. MCP is a client-server protocol that enables a single agent to discover, authenticate with, and invoke external tools and data sources in a standardized manner. The protocol establishes clear interfaces for capability negotiation, parameter validation, and result handling, allowing agents to safely integrate with tools, resources, and prompts. MCP’s architecture is a hub-and-spoke model where the agent coordinates connections to multiple tool providers.
Agent-to-Agent (A2A) Protocol: Multi-Agent Coordination. A2A protocol addresses the complementary challenge of inter-agent communication and delegation in multi-agent systems [5]. Unlike MCP’s focus on individual agent-tool integration, A2A specifically enables horizontal coordination between multiple autonomous agents. The protocol provides mechanisms for agents to discover peer capabilities, negotiate task assignments, and coordinate execution across distributed environments. A2A operates as an open protocol built upon established standards such as HTTP, Server-Sent Events (SSE), and JSON-RPC, which allows for seamless integration with existing enterprise infrastructures and promotes interoperability across various agentic AI.
Central to the A2A protocol is the Agent Card, a metadata structure that encapsulates an agent’s capabilities, authentication requirements, and communication endpoints. Through Agent Cards, agents advertise functionalities and establish trust with other agents, enabling dynamic collaboration. The protocol’s design emphasizes agentic autonomy, allowing agents to collaborate in their natural, unstructured modalities without shared memory or tools. This approach supports multi-agent scenarios, where agents delegate tasks to specialized counterparts, thereby optimizing overall system performance.
III Challenges in Compositional Reasoning of Agentic AI Systems
@@
The integration of the capabilities provided by the MCP and A2A protocols within Agentic AI systems creates a composition of multiple components. These components operate on fundamentally different abstractions and assume distinct trust models [6]. For instance, MCP treats external tools as trusted resources with well-defined interfaces, whereas A2A handles potentially untrusted peer agents with varying capabilities and reliability guarantees.
The resulting lack of a unified semantic model creates two primary classes of risk: ($a$) architectural misalignment and ($b$) exploitable coordination issues.
@@
Architectural Misalignment and Semantic Gap. The core difficulty raises from the lack of a unified semantic layer between MCP’s tool-centric model and A2A’s agent-centric model. This misalignment makes it difficult to reason about correctness properties, maintain state consistency, and verify security guarantees when both protocols are used simultaneously within the same system [11]. The complexity is further amplified by the dynamic nature of agents, which may exhibit behaviors that violate protocol assumptions [33]. This creates a lack of formal guarantees in two areas:
@@
Coordination Issues. The absence of tools to validate the correctness of cross-protocol interactions creates design flaws that an adversary can exploit.
  • Circular Delegation Loops (Deadlocks): This issue occurs when tasks pass indefinitely between agents. For example, Agent A delegates to Agent B (A2A), which in turn delegates back to Agent A (A2A), creating a cycle that halts progress.
  • Privilege Escalation: This issue occurs when a malicious agent exploits delegation chains to gain unauthorized access to tools. For instance, Agent A, which lacks direct access to a sensitive MCP tool, delegates an innocent-sounding task to Agent B (A2A), which possesses the necessary MCP credentials, thereby using Agent B as a proxy for unauthorized access.
@@
Figure 1: Overview of the host agent model with an Agentic AI system. The model mediates user interactions and coordinates task execution across a heterogeneous set of external entities, specifically MCP [4] and A2A [5] servers. Internally, the HA employs a modular design comprising the Host Agent Core ( A), the Orchestrator ( C), and the Communication Layer ( D). The execution lifecycle proceeds as follows: A processes the user prompt and resolves intent; B enables dynamic discovery of external entities; A, B, and C collaborate on DAG-based sub-task planning. Execution involves secure task invocation managed by D, followed by aggregation and resolution of results by C before the final response returns to the user. This architecture supports scalable, interpretable, and compositional collaboration among autonomous AI agents and tools ( E).
IV Modeling the Agentic AI Systems
We abstract the formal model for the Host Agent (HA) to provide a unified view of an agentic AI system. The HA receives natural language requests from users, decomposes them into structured subtasks, assigns those subtasks to AI agents or other external entities, and aggregates the results. It also oversees execution, monitors progress, and manages exceptions or recovery procedures.
@@ IV Modeling the Agentic AI Systems
We then complement the host agent model with the task lifecycle model, which specifies the states and transitions a sub-task undergoes from its origin to completion. This model offers a fine-grained view of task management and captures state transitions and dependencies. Our deliberate separation of the host agent and task lifecycle models serves two main purposes. First, it establishes clear abstraction boundaries to delineate responsibilities; the HA handles orchestration and intent resolution, while the task model manages the execution details of individual sub-tasks. Second, this separation allows integration of new sub-task states to the task model without requiring fundamental changes to the host agent model.
@@
We derive these models by synthesizing the operational structures, specifications, and reference implementations of the unified protocols (Anthropic’s MCP [4] and Google’s A2A [5]). This methodology incorporates insights from modular orchestration and dynamic task planning practices for Agentic AI systems [21].
@@
The architectural principles of the HA are inspired by modern orchestration frameworks such as LangGraph [34], n8n [35], and OpenAI’s AgentKit [36]. However, these frameworks typically execute predefined, graph-based workflows. Our model advances this paradigm by empowering the HA to dynamically generate and adapt execution plans in response to prompted user requests. This enables autonomous coordination, moving beyond the constraints of static flowcharts. Similarly, our task lifecycle model generalizes state management concepts found in protocols such as A2A. We extend this structure to support more fluid and emergent task progressions, where state transitions are determined by runtime conditions and autonomous agent decisions (e.g., failures and fallbacks).
@@
We leverage these models to abstract key operational structures of Agentic AI systems. This abstraction provides a foundation for property identification aimed at safe system design through the formalization of task delegation, inter-agent coordination, and execution oversight, detailed in Section V.
IV-A Constructing the Host Agent Model
HA is responsible for interacting with users, understanding and reasoning through queries, and selecting external entities, e.g., tools and AI agents.
Figure 1 illustrates its abstracted model in an Agentic system. The primary function of HA is to manage the lifecycle of complex tasks, from initial reception to completion. This includes initiating a strategic context request, a process where the HA uses internal state, historical records, and external data to gather situational information for accurate task decomposition, optimal tool selection, and multi-agent coordination.

### v2

@@ Abstract
Agentic AI systems, which leverage multiple autonomous agents and large language models (LLMs), are increasingly used to address complex, multi-step tasks. The safety, security, and functionality of these systems are critical, especially in high-stakes applications. However, the current ecosystem of inter-agent communication is fragmented, with protocols such as the Model Context Protocol (MCP) for tool access and the Agent-to-Agent (A2A) protocol for coordination being analyzed in isolation. This fragmentation creates a semantic gap that prevents the rigorous analysis of system properties and introduces risks such as architectural misalignment and exploitable coordination issues. To address these challenges, we introduce a modeling framework for agentic AI systems composed of two central models: (1) the host agent model formalizes the top-level entity that interacts with the user, decomposes tasks, and orchestrates their execution by leveraging external agents and tools; (2) the task lifecycle model details the states and transitions of individual sub-tasks from creation to completion, providing a fine-grained view of task management and error handling. Together, these models provide a unified semantic framework for reasoning about the behavior of multi-AI agent systems. Grounded in this framework, we define 16 properties for the host agent and 14 for the task lifecycle, categorized into liveness, safety, completeness, and fairness. Expressed in temporal logic, these properties enable formal verification of system behavior, detection of coordination edge cases, and prevention of deadlocks and security vulnerabilities. Through this effort, we introduce the first rigorously grounded, domain-agnostic framework for the analysis, design, and deployment of correct, reliable, and robust agentic AI systems.
1 Introduction
@@
Agentic AI systems are becoming central to the creation of advanced AI. These systems address complex, multi-step tasks that surpass single-agent capabilities by utilizing multiple autonomous agents. Agents in these systems employ Large Language Models (LLMs) as their primary reasoning engine, which enables them to perform task planning, delegation, and complex decision-making (Park et al., 2023; Nascimento et al., 2023; Zhang et al., 2024). These heterogeneous agents (e.g., agents based on language, vision, robotics, and symbolic planning) dynamically communicate, delegate, and collaborate to complete tasks.
@@
To illustrate, we consider an automated financial planner. A host agent accepts a natural language request from a user to “create a budget and invest $\mathdollar 1,000$”. It decomposes this into subtasks: a data agent gathers market data, a planning agent calculates risk, and a transaction agent executes the final investment. These specialized agents interact and exchange data to accomplish the multi-step task. The operation of such systems relies on standardized communication protocols, such as Model Context Protocol (MCP) (Anthropic, 2024), which addresses vertical tool access for a single agent, and Agent-to-Agent (A2A) (Google Cloud, 2025) protocol, which addresses horizontal agent coordination via inter-agent delegation.
@@
Turning back to the financial planner example, the agentic system must leverage both protocols. The host agent uses A2A to delegate the subtask of gathering market data to the data agent (horizontal agent coordination). The transaction agent relies on MCP to invoke an external banking API tool to execute the final investment (vertical tool access). The host agent’s ability to manage and sequence both A2A and MCP interactions across various agents enables the multi-step task to succeed.
@@
Formal or empirical assessment of the safety, security, and functionality of multi-AI agent systems is critical as these systems transition to open-ended, high-stakes applications (de Witt, 2025; Han et al., 2025). Here, safety involves preventing the agent’s actions from causing unintended or harmful physical or real-world consequences, such as an incorrect trade that causes a substantial loss in the financial planner example. Functionality ensures that the system accurately and reliably completes the specified task, such as verifying that the financial planner calculates risk correctly and executes the investment as requested. Safety and functionality issues may arise naturally, e.g., due to design flaws or code errors, whereas security considers adversarial manipulation and unauthorized access that can violate these requirements. For instance, a malicious external agent could use a prompt injection attack (Liu et al., 2024b) to force the transaction agent into transferring funds to an unauthorized account, a clear violation of both security and safety policies.
@@
Recent work directly or indirectly studies these issues by examining agents in specific settings, such as adversarial agent behavior (Huang et al., 2025; He et al., 2025; Peigné et al., 2025) and protocol misuse in task delegation (Motwani et al., 2024). Others focus on the compromise or malfunction of a single agent that propagates errors, escalates failures, or exposes vulnerabilities across an entire system (Tian et al., 2024; Lee and Tiwari, 2025; Zhang et al., 2021; Hedin and Moradian, 2015). Further studies examine the risk of using combinations of safe models for misuse (Jones et al., 2025), the security challenges inherent in tool usage (Ye et al., 2024), and ensuring secure delegation and preventing protocol misuse in adversarial settings (Habler et al., 2025; Li and Xie, 2025).
@@
The current ecosystem of inter-agent communication and coordination is fragmented and lacks a cohesive formal foundation. Existing protocols (e.g., MCP and A2A) are analyzed in isolation, which creates a semantic gap that impedes rigorous reasoning about safety, security, and functional properties when both are used simultaneously in complex workflows. This absence of a unified framework, a need that is also recognized by multi-stakeholder initiatives (Neelou and others, 2025), leaves deployments vulnerable to emergent coordination failures, including deadlocks and privilege escalation, that cannot be adequately verified or mitigated. Achieving verifiable correctness in these systems requires a rigorous, unified modeling framework that can capture the execution behavior of diverse agents across heterogeneous protocols.
To address these challenges, we introduce a two-model framework for agentic AI systems:
@@
    The Host Agent Model formalizes the top-level entity that interacts with the user. It is responsible for accepting user tasks, decomposing them into structured subtasks, and orchestrating execution via external entities (agents via A2A, tools via MCP). It acts as both a controller and a monitor, ensuring safe delegation and state consistency throughout the task lifecycle.
@@
    The Task Lifecycle Model formalizes the dynamic structure of task management, detailing states and transitions a sub-task undergoes from its origin to its completion or failure. It provides the fine-grained execution logic necessary for dependency management and resilient error handling.
These models are the foundational structure that enables formal verification. They capture discrete states and transitions of the entire coordination process and provide the necessary semantic framework (state space) to express the formal safety, security, and functional requirements. This unified perspective is required to rigorously reason about execution behavior, uncover coordination edge cases, and verify properties critical to real-world deployments.
We categorize the formal properties defined within this framework as liveness, safety, completeness, and fairness, which are essential for verifiable system assurance. For instance, a safety property guarantees that the transaction agent will not execute an investment before the planning agent has successfully calculated the risk, preventing inconsistent state or premature execution. Conversely, a liveness property ensures that once a user submits a request, a final response is eventually returned. This prevents a task from entering a permanent deadlock or starvation state.
Contributions. In summary, we make the following contributions:
  • We introduce a host agent model, providing a rigorous and domain-agnostic foundation for the systematic analysis of multi-agent systems. We formalize the top-level agent that interacts with users, decomposes tasks, and orchestrates their execution by leveraging external agents and tools.
@@
2 Background
Agentic AI systems. AI agents are autonomous entities that perceive, reason, and act to achieve goals (Russell and Norvig, 2009). While early systems were symbolic (Nilsson, 1998), modern agents use LLMs to reason, plan, and execute tool-use tasks (Wang et al., 2024; Durante et al., 2024; Xi et al., 2025). However, complex multi-faceted tasks often exceed single-agent capabilities. Agentic AI systems address this by forming collaborative networks (Stone and Veloso, 2000; Dorri et al., 2018), where specialized agents interact to solve problems beyond individual scope through dynamic task decomposition and allocation (Tampuu et al., 2017; Li et al., 2023; Park et al., 2023).
@@ 2 Background
Agent Communication and Coordination. Agentic AI systems require standardized frameworks for individual agent capabilities and multi-agent coordination. These frameworks comprise two complementary aspects: protocols for agents to interact with external resources and tools, and protocols for inter-agent communication. For individual agents, the challenge is ensuring secure, reliable access to external computational resources and tools. Agents must discover capabilities, authenticate services, validate parameters, and handle errors robustly. For multi-agent coordination, protocols define how agents discover peer capabilities, negotiate task assignments, exchange information, and maintain consistency across distributed execution contexts.
The MCP (Anthropic, 2024) and A2A (Google Cloud, 2025) protocols illustrate the integration challenges that arise from combining individual agent capabilities with multi-agent coordination. MCP standardizes agent-tool integration (Anthropic, 2024). Operating as a client-server protocol, it enables agents to discover, authenticate and execute commands across external data sources and toolsets. This hub-and-spoke model ensures agents can retrieve context and trigger actions without requiring unique code for every provider.
While MCP facilitates tool usage, A2A enables horizontal coordination between autonomous agents (Google Cloud, 2025). Built on HTTP and JSON-RPC, A2A allows agents to discover peer capabilities and negotiate task assignments across distributed environments. Central to A2A is the Agent Card, metadata describing an agent’s capabilities and endpoints. This allows agents to establish trust and delegate tasks dynamically without shared memory.
3 Challenges in Compositional Reasoning of Agentic AI Systems
@@
Integrating MCP and A2A within Agentic AI systems creates a composition of multiple components. These operate on different abstractions and assume distinct trust models (de Witt, 2025). For instance, MCP treats external tools as trusted resources with well-defined interfaces, whereas A2A handles potentially untrusted peer agents with varying capabilities and reliability guarantees.
@@
Architectural Misalignment and Semantic Gap. The core difficulty arises from the lack of a unified semantic layer between MCP’s tool-centric model and A2A’s agent-centric model. This misalignment makes it difficult to reason about correctness properties, maintain state consistency, and verify security guarantees when both protocols are used simultaneously within the same system (Peigné et al., 2025). The complexity is further amplified by the dynamic nature of agents, which may exhibit behaviors that violate protocol assumptions (Hammond et al., 2025). This creates a lack of formal guarantees in two areas:
@@
Coordination Issues. The absence of tools to validate the correctness of cross-protocol interactions creates design flaws, such as functional collapse and security breaches, that an adversary can exploit. First, circular delegation and deadlock emerge when tasks pass indefinitely between agents. For example, Agent A delegates to Agent B (A2A), which in turn delegates back to Agent A (A2A), creating a cycle that halts progress. Second, privilege escalation occurs when a malicious agent exploits delegation chains to gain unauthorized access to tools. For instance, Agent A, which lacks direct access to a sensitive MCP tool, delegates an innocent-sounding task to Agent B (A2A), which possesses the necessary MCP credentials, thereby using Agent B as a proxy for unauthorized access.
@@
4 Modeling the Agentic AI Systems
@@
We abstract the formal model for the Host Agent (HA) to provide a unified view of an Agentic AI system. The HA receives natural language requests from users, decomposes them into structured subtasks, assigns those subtasks to AI agents or other external entities, and aggregates the results. It also oversees execution, monitors progress, and manages exceptions or recovery procedures. We then complement the host agent model with the task lifecycle model, which specifies the states and transitions a sub-task undergoes from its origin to completion. Our deliberate separation of the host agent and task lifecycle models serves two main purposes. First, it establishes clear abstraction boundaries to delineate responsibilities; the HA handles orchestration and intent resolution, while the task model manages the execution details of individual sub-tasks. Second, it allows integration of new sub-task states to the task model without requiring fundamental changes to the HA model.
@@
We derive these models by synthesizing the operational structures, specifications, and reference implementations of the unified protocols (Anthropic’s MCP and Google’s A2A). This incorporates insights from modular orchestration and dynamic task planning practices for Agentic AI systems.
@@
Recent orchestration frameworks such as LangGraph (LangChain, 2025), n8n (n8n, 2025), and OpenAI’s AgentKit (OpenAI, 2025) inspire the architectural principles of the HA. However, these frameworks typically execute predefined, graph-based workflows. The proposed model extends this paradigm. The HA creates and adjusts execution plans dynamically in response to user requests. This approach permits autonomous coordination and avoids the constraints of static flowcharts. Similarly, the task lifecycle model generalizes state management concepts from protocols such as A2A. This structure supports flexible task progressions. In this framework, runtime conditions and autonomous agent decisions (e.g., failures and fallbacks) determine state transitions.
@@
We leverage these models to abstract key operational structures of Agentic AI. This abstraction provides a foundation for property identification aimed at safe system design through the formalization of task delegation, inter-agent coordination, and execution oversight, detailed in Section 5.
Figure 1: The Host Agent architecture.
4.1 Constructing the Host Agent Model
The Host Agent interprets users’ queries, selects external entities, and manages task lifecycles (Figure 1). It uses internal state, historical records, and external data to gather context for task decomposition and agent coordination.